# Localization landscape of optical waves in multifractal photonic membranes

### 26 Jan 2024 v1

@@ 1 Introduction
In this article, we propose a general method to calculate the landscape function of classical optical waves in multifractal structures based on the Helmholtz equation. First, we formulate the eigenproblem based on Helmholtz equation of both one-dimensional (1D) and two-dimensional (2D) multifractal systems, which are analogous to the Hamiltonian operator of the Schrödinger equation. Next, we discuss the generation of the investigated multifractal scattering structures and then we calculate the associated localization landscapes by solving the Dirichlet problem for the classical optical potential. Based on the obtained Helmholtz landscapes, we design and fabricate nanoperforated multifractal membranes in SiN and investigate experimentally their optical resonances using leaky-mode imaging spectroscopy.
@@ 2 Generation of multifractal scattering structures
Figure 1: Constructed multifractal potential by (a) the 1D multiplicative cascade method and (b)the MMAR method. Also shown in (c) and (d) are landscape functions corresponding to (a) and (b), calculated for $\overline{d}_{1}/\lambda=0.4$.
@@ 3 The Helmholtz localization landscape of one-dimensional multifractals
For the 1D geometry, the complex amplitude of an harmonic optical wave $\psi(x)$ is governed by Helmholtz equation:
$\left[\Delta+k^{2}\right]\psi(x)=0$ (2)
where $k^{2}=\overline{k}^{2}-V_{s}(x)$ is the fluctuating wavenumber (around the average wavenumber $\overline{k}$) related to the averaged permittivity $\overline{\epsilon_{r}}$ of the medium by $\overline{k}^{2}=\overline{\epsilon_{r}}k_{0}^{2}$ and $\Delta$ denotes the Laplacian operator. Here $k_{0}=\omega/c$ is the free-space wavenumber and $V_{s}(x)$ is the spatially-dependent optical scattering potential, which is proportional to the local fluctuations of the permittivity of the scattering medium according to [6]:
$V_{s}(x)=k_{0}^{2}\left[\epsilon_{r}(x)-1\right].$ (3)
We can rewrite the Helmholtz equation above for the non-homogeneous medium as the equivalent eigenvalue problem:
$\left[-\Delta+V_{s}(x)\right]\psi(x)=\overline{\epsilon_{r}}k_{0}^{2}\psi(x).$ (4)
that has the same structure of the Schrödinger equation. Its solutions depend on the spatially disordered optical potential $V_{s}(x)$ [10].
The Hamiltonian of the system, defined by $H=-\Delta+V_{s}(x)$, allows us to determine the localization landscape function $u(x)$ that encodes the fundamental information on the localized eigenmodes [15, 16, 17] through the solution of the associated Dirichlet problem:
@@
$-t\left[u_{i+1}+u_{i-1}-2u_{i}\right]+\left(V_{i}-2t\right)u_{i}=1$ (6)
To make sure that the landscape function $u_{i}$ is positive everywhere, the optical scattering potential term should satisfy the additional condition that $V_{i}-2t\geq 0$ [16]. Note that given the fact that the potential $V_{s}$ is a bounded non-negative function, one can define the so-called "effective potential" $W=1/u$. The profiles of the effective potential $W$ are often much more regular than the physical potential $V$ and display clear structures of walls and wells that identify regions of localization. A rigorous connection can be made between the exponential decay of eigenfunctions and the effective potential $W$ [17]:
@@
In the calculation of the landscape function using Eq. 6, we considered the hopping rate to be $t=1$. Our potentials take piecewise constant values of $\epsilon_{r}$ and $0.2\epsilon_{r}$ based on whether a particle is present at a given point along the chain or is missing. The multifractal potentials generated by the two described methods are shown in Figure 1(a,b). We selected the permittivity of the scattering medium to be $\epsilon_{r}=10.5$ [30], and the incident wavelength $\lambda$ is chosen to satisfy $\overline{d_{1}}/\lambda=0.4$ where $\overline{d_{1}}$ is the average inter-particle separation along the chain. The selected $\overline{d_{1}}/\lambda$ is typical for the formation of bandgaps in photonic crystal structures [22].
The landscape functions $u(x)$ calculated by Eq. 5 are plotted in Figure 1(c,d). To demonstrate the accuracy of the localized mode predictions from the obtained landscape function, we directly calculated the eigenmodes from Eq. 4 of both structures and plotted some representative modes in Figure 2(a,b). We observed the coexistence of both localized modes with exponential decay in space and modes with large intensity fluctuations and reduced localization behavior as well. Moreover, in Figures 2(c,d) we compared the localized mode positions predicted by the peaks of the landscape function $x_{\mathrm{pred}}$ and the actual positions $x_{\mathrm{actual}}$ of the modes obtained numerically, selecting a wide range of modes that are spatially distributed along the entire line segment. Our data demonstrate almost perfect correlation between the predicted and the actual positions of the localized modes based on the computed landscape of the Helmholtz operator.
We now apply the proposed landscape approach to more challenging 2D multifractal systems generated by the multiplicative cascade method as in reference [20]. Therefore, we consider the natural generalization of Eq. 2:
$H\psi(\mathbf{r})=\overline{\epsilon_{r}}k_{0}^{2}\psi(\mathbf{r}).$ (8)
Figure 4: (a-i) Typical eigenmodes obtained at $\overline{d}_{1}/\lambda=0.55$ for the investigated monofractal structure.
where the 2D Hamiltonian operator is defined as $H=-\Delta+V_{s}(\mathbf{r})$. The associated localization landscape function $u(\mathbf{r})$ is obtained by solving the Dirichlet problem:
$Hu(\mathbf{r})=1.$ (9)
where, similarly to the 1D case, $V_{s}(\mathbf{r})\geq 0$ everywhere in $\mathbf{r}=(x,y)$.
Figure 5: (a-i) Typical eigenmodes obtained at $\overline{d}_{1}/\lambda=0.82$ for the investigated multifractal structure.
@@ 4 The Helmholtz localization landscape of two-dimensional multifractals
The landscape functions are calculated by solving Eq. 9. We set up the 2D Laplacian operator with Dirichlet boundary conditions on a domain $\Omega$ and discretized the operators using high-order finite elements using triangular and tetrahedral elements. In order to obtain more accurate results, we oversampled the domain $\Omega$ where $2\times 2$ elements per constant piece of the optical potential. We need to emphasize that the optical potentials just refers to the coordinates of scatterers and do not take into account the radii of each scatterer. That is to say we treat the scattering medium to be a point distribution, which is not the case for the fabricated nanohole membranes. However, such approximation does not affect the accuracy of the capability of the method to predict eigenmodes because the fabricated nanoholes have very small radii comparing to the multiscale nature of the scatterer separation. We developed MATLAB versions that implement the localization landscape theory. The multifractal potential term is added to the discrete Laplacian operator to obtain the discrete Helmholtz operator $H$. Our numerical simulations were performed using the Boston University’s Shared Computing Cluster (SCC) [31].
@@
In our simulations, we selected the material permittivity to be $\epsilon_{r}=4$, corresponding to the value for the silicon nitride (SiN) material used in the fabrication of our devices. We selected $\overline{d}_{1}/\lambda=0.55$ and $\overline{d}_{1}/\lambda=0.82$ to compute the landscape and the eigenmodes of monofractal and multifractal systems respectively. The selected $\overline{d}_{1}/\lambda$ lies in the typical range for a photonic crystal structure where bandgaps are likely to form [30, 22] and optical modes with high-quality factors appear at the band-edges [30, 32, 33]. We emphasize that the range of the selected spectral parameter $\overline{d}_{1}/\lambda$ used in the simulations overlaps with the one of the fabricated structures that are investigated in the experimental section of this paper.
@@
The landscape function $u(\mathbf{r})$ of both monofractal and multifractal structures are plotted in Figures 3(c,d). One can see that the landscape of the monofractal potential features intense maxima localized within small clusters of adjacent scattering particles distributed across structure and with spatial distributions that closely resemble its geometrical fractal support. These are the regions in which the localization landscape predicts the existence of highly confined resonant modes localized by proximity effects occurring among closely coupled particles, which is the typical behavior observed for small fractal aggregates of dipolar particles [34]. On the other hand, the landscape of the multifractal potential displays more extended localization regions unevenly distributed across the structure, consistent with its non-homogeneous multiscale nature [20]. We plotted several representative modes for both the monofractal and multifractal structures in Figure 4 and Figure 5, respectively. As predicted by the general nature of the localization landscape, we observe for the monofractal cluster strongly localized modes due to local dipole coupling effects, which are known to occur in self-similar arrays where they have been succesfully exploited to dramatically increase the cross sections of nonlinear and Raman processes [34]. However, the situation is very different in the investigated multifractal structure where we observe modes with strongly fluctuating amplitude distributions and a broad range of localization lengths. In the next sections, we discuss the fabrication and characterization of dielectric membrane structures designed based on the localization landscape theory and we demonstrate experimentally the formation of their characteristic fractal/multifractal modes in good qualitative agreement with the landscape predictions.
@@ 5.1 Silicon nitride thin film growth and optical characterization
In this section we discuss the optical characterization of highly transparent silicon nitride thin films grown by reactive magnetron sputtering. Silicon nitride has become a well established material platform for integrated Si photonics where devices such as ultralow-loss waveguides [35], high-Q resonators, and integrated structures for telecommunications, sensing, and metrology have been demonstrated [36]. In this work, silicon nitride thin films were grown atop of silicon and fused silica substrates via reactive RF magnetron sputtering using a Denton Discovery 18 sputtering system with a base pressure of $2\times 10^{-7}$orr and a substrate temperature held at $300^{\circ}$. We used a 99.99% purity 3-in silicon target and sputtered the thin films in a 2:1 argon-nitrogen environment at 2.5mTorr deposition pressure and 200W of RF power resulting in a deposition rate of $\approx 5\,\mathrm{nm/min}$. All substrates were washed in a hot piranha solution and plasma ashed in an oxygen environment prior to deposition. We characterize the optical constants of the fabricated SiN thin films from the ultra-violet (UV) to the near infrared (NIR) via broadband variable angle spectroscopic ellipsometry (VASE) and normal incidence transmission measurements. In spectroscopic ellipsometry, a parameterized oscillator model is introduced to capture the relevant aspects of a materials optical dispersion as a function of wavelength. The oscillator model is then used to interpret the measured data and extract the optical constants of materials via accurate fitting procedures [37]. VASE measurements at three angles (${}^{\circ}$${}^{\circ}$${}^{\circ}$) were performed on SiN thin films grown atop of silicon substrates along with normal incidence transmission measurements of the same films grown atop of fused silica substrates. Figure 6 summarizes our work on characterizing the optical properties of the fabricated SiN thin films. Panels (a-b) show the measured and fitted ellipsometric parameters $\Psi$ and $\Delta$ corresponding to the magnitude and phase of the ratio of the complex s- and p-polarized Fresnel reflection coefficients, respectively, at each angle along with each respective fit. The complex Fresnel reflection coefficients are related to the complex refractive index and thickness of the material through parameterized oscillator models. In particular, we employ a single parameterized Tauc-Lorentz (TL) oscillator model which has been used to extract the optical constants of amorphous materials in excellent agreement with experimental measurements [38]. We find that this modeling method gives nearly ideal agreement with our measurements over the entire UV-VIS-NIR region investigated with a mean square error (MSE) of less than 10. In Figure 6(c) we report the extracted optical constants of the fabricated 350nm thick SiN thin film as a function of wavelength. Furthermore, we performed normal incidence transmission measurements on the SiN thin films grown atop of fused silica substrates and independently test our TL model by plotting the expected transmission. We report our measurements and fit in fig 6(d) finding the TL model is in excellent agreement with measurement over the entire wavelength range investigated. The transmission spectra, along with the extracted optical constants of the fabricated material, demonstrates exceptionally high transparency with an almost constant refractive index of $2$. Consistent with literature, we attribute these optical properties to the high RF power and substrate temperature during deposition[39]. In the next section we discuss the fabrication of nanohole perforated membranes with multifractal geometries.

### 05 Mar 2024 v2

@@ 1 Introduction
In this article, we propose a general method to calculate the landscape function of classical optical waves in multifractal structures based on the Helmholtz equation. First, we formulate and solve an eigenproblem analogous to the Hamiltonian operator of the Schrödinger equation and validate our results by considering the well-known problems of one-dimensional waveguide structures. We then apply the developed approach to both one-dimensional (1D) and two-dimensional (2D) multifractal systems. Next, we discuss the generation of the investigated multifractal scattering structures and then we calculate the associated localization landscapes by solving the Dirichlet problem for the classical optical potential. Based on the obtained Helmholtz localization landscapes, we design and fabricate nanoperforated multifractal membranes in SiN and investigate experimentally their optical resonances using leaky-mode imaging spectroscopy.
@@ 3 The Helmholtz localization landscape of one-dimensional multifractals
Our goal is to apply the localization landscape theory (LLT) to photonic problems governed by the scalar Helmholtz equation in a scattering dielectric medium with arbitrary geometry. We start from the Helmholtz equation in a non-homogeneous medium:
$\left[\Delta+k_{0}^{2}\epsilon(x)\right]\psi=0$ (2)
where $k_{0}$ is the free-space wavenumber and $\epsilon(x)$ is the spatially varying permittivity of the medium. We then define the variation from the background medium in terms of the optical scattering potential $V_{s}(x)$ as follows [6]:
$k^{2}=k_{0}^{2}\epsilon(x)=k_{b}^{2}-V_{s}(x)$ (3)
where $k^{2}_{b}=k_{0}^{2}\epsilon_{b}$, and $\epsilon_{b}$ is the permittivity of the background medium (we use $\epsilon_{b}=1$). From the expression above we have:
$V_{s}(x)=k_{0}^{2}\left[\epsilon_{b}-\epsilon(x)\right].$ (4)
We can now formally express the Helmholtz equation as the equivalent Schrödinger equation:
$\left[-\Delta+V_{s}(x)\right]\psi=k_{0}^{2}\epsilon_{b}\psi.$ (5)
However, the potential $V_{s}(x)$ can become negative when the permittivity of the scattering particles is larger than the one of the background (host medium). Therefore, in order to apply the LLT we implement the following transformation:
$V_{s}(x)\rightarrow V^{\prime}(x)\equiv V_{s}(x)+k_{0}^{2}\left[\epsilon_{\mathrm{max}}-\epsilon_{b}\right]=k_{0}^{2}\left[\epsilon_{\mathrm{max}}-\epsilon(x)\right].$ (6)
were $\epsilon_{\mathrm{max}}$ is the maximum value of the permittivity of the scattering medium. The shift introduced above renders the potential non-negative, which enables the application of the LLT to general optical wave problems without affecting their physical solutions. We now introduce the Helmholtz landscape equation based on the Dirichlet solution of the equation:
@@
where we defined the optical Hamiltonian $H=-\Delta+V^{\prime}(x)$. Equation 7 allows us to investigate the localization landscape function $u(x)$ that encodes fundamental information on the localized eigenmodes [15, 16, 17].
Figure 1: Tight-binding optical potential of (a) single waveguide and (b) coupled waveguide system. (c,d) Spatial profiles of the lowest 3 optical modes from the solution of Eq. 5, along with the calculated landscape function $u$ (doted lines) computed from Eq. 7.
@@
$-t\frac{\left[u_{i+1}+u_{i-1}-2u_{i}\right]}{\delta^{2}}+\left(V_{i}-\frac{2t}{\delta^{2}}\right)u_{i}=1$ (8)
where $\delta$ is the discretization step of the coordinate $x$-axis. To ensure that the spectrum of the possible eigenenergies is positive everywhere, the optical scattering potential term should satisfy the additional condition that $V_{i}-2t/\delta^{2}\geq 0$ [16]. Note that since the potential $V^{\prime}$ is a bounded non-negative function, one can also define the so-called "effective potential" $W=1/u$. The profiles of the effective potential $W$ are often more regular than the optical potential and display clear structures of walls and wells that serve to identify the regions of localization of the eigenmodes. A rigorous connection has been made between the exponential decay of eigenfunctions and the effective potential $W$ [17]:
@@
Figure 2: Constructed multifractal potential by (a) the 1D multiplicative cascade method and (b)the MMAR method. Also shown in (c) and (d) are square root of landscape functions, normalized to their maxima corresponding to (a) and (b), calculated for $\overline{d}_{1}/\lambda=0.4$. First five modes of each potential are overlayed on top of the landscape function, normalized to their intensities respectively.
In order to demonstrate the predictive power of the Helmholtz landscape introduced above, we consider first the simple case of a three-layer asymmetric waveguide structure with a dielectric permittivity in the core region $\epsilon_{c}=\epsilon_{max}=12.25$ and with cladding permittivities $\epsilon_{1}=1$ and $\epsilon_{2}=2.10$. We use a wavelength $\lambda=1.55\mu$m and display the tight-binding optical potential from Eq. 6 and Eq. 8 in Figure 1(a). In addition, we considered a system of two coupled dielectric waveguides with the optical potential shown in Figure 1(b). These are canonical eigenproblems in photonics whose solutions can easily be obtained with traditional methods [30]. Here, by computing the eigensolutions of the corresponding Schrödinger equation 5, we plot the first 3 optical modes of the two investigated systems in Figures 1(c,d), respectively. The dashed lines indicate the corresponding Helmholtz landscapes obtained by solving equation 7. In all the calculations we set the hopping rate to be $t=1$, without loss of generality. The results plotted in Figure 1 indicate that the Helmholtz landscape accurately predicts the localization regions of the optical modes computed from the tight-binding solution of the equivalent Schrödinger problem.
We then apply the developed approach to the more challenging cases of fractal and multifractal optical potentials that take piecewise constant values equal to $\epsilon_{r}$ and 1 based on whether a particle is present at a given point of the chain or is missing. The multifractal potentials generated by 1D multiplicative cascade method and by the MMAR method are shown in Figure 2(a,b), respectively. We selected the permittivity of the scattering medium to be $\epsilon_{r}=10.5$ [31], and the incident wavelength $\lambda$ is chosen to satisfy $\overline{d_{1}}/\lambda=0.4$ where $\overline{d_{1}}$ is the average inter-particle separation along the chain. The selected $\overline{d_{1}}/\lambda$ is typical for the formation of bandgaps in photonic crystal structures [22].
The landscape functions $u(x)$ calculated by Eq. 7 are plotted in Figure 2(c,d). To demonstrate the accuracy of the localized mode predictions from the obtained landscape function, we directly calculated the eigenmodes from Eq. 5 of both structures and plotted some representative modes in Figure 3(a,b). Particular care has been taken to ensure that the calculated eigenvectors correspond to the eigenvalue of Eq. 5, determined by the free-space background medium, which are narrowly distributed around $k_{0}^{2}$ (within a $5\%$ dispersion value). We observed the coexistence of both localized modes with exponential decay in space and modes with large intensity fluctuations and reduced localization behavior as well. Moreover, in Figures 3(c,d) we compared the localized mode positions predicted by the peaks of the landscape function $x_{\mathrm{pred}}$ and the actual positions $x_{\mathrm{actual}}$ of the modes obtained numerically, selecting a wide range of modes that are spatially distributed along the entire line segment. Our data demonstrate almost perfect correlation between the predicted and the actual positions of the localized modes based on the computed landscape of the Helmholtz operator.
We now apply the Helmholtz landscape approach to more challenging 2D multifractal systems generated by the multiplicative cascade method as in reference [20]. Therefore, we consider the natural generalization of Eq. 2:
$H\psi(\mathbf{r})=k_{0}^{2}\epsilon_{b}\psi(\mathbf{r}).$ (10)
where the 2D Hamiltonian operator is defined as $H=-\Delta+V^{\prime}(\mathbf{r})$. The associated localization landscape function $u(\mathbf{r})$ is obtained by solving the Dirichlet problem:
$Hu(\mathbf{r})=1.$ (11)
where, similarly to the 1D case, $V^{\prime}(\mathbf{r})\geq 0$ everywhere in $\mathbf{r}=(x,y)$.
@@ 4 The Helmholtz localization landscape of two-dimensional multifractals
The landscape functions are calculated by solving Eq. 11. We wish to setup the Laplacian operator with homogeneous Dirichlet boundary conditions on an open bounded domain $\Omega\in\mathbb{R}^{2}$ with boundary $\partial\Omega$. We discretize using finite elements with linear basis functions. Given a bounded, symmetric bilinear form111in our case, $a(u,v)=\int_{\Omega}\nabla u\cdot\nabla v$ $a(u,v)$ that is coercive on $H^{1}_{0}(\Omega)$, we want to find $u\in H^{1}_{0}(\Omega)$ such that $u$ satisfies
$a(u,v)=0,\forall v\in H^{1}_{0}(\Omega),$
@@
where $H^{1}_{0}(\Omega)\subset L_{2}(\Omega)$, denotes the subspace of functions with square integrable derivatives that vanish on the boundary. This problem is known to have a unique solution $u^{*}$ [32]. For the numerical comparisons in this work we consider domains that are the image of a square under a diffeomorphism, i.e., a smooth mapping from the reference domain $S:=[0,1]^{2}$ to the physical domain $\Omega$. Quadrilateral finite element meshes and tensorized nodal basis function based on Legende-Gauss-Lobotto (LGL) points are used. We use isoparametric elements to approximate the geometry of $\Omega$, i.e., on each element the geometry diffeomorphism is approximated using the same basis functions as the finite element approximation. The Jacobians for this transformation are computed at every quadrature point, and Gauss quadrature is used to numerically approximate integrals. We restrict our comparisons to uniformly refined conforming meshes and our implementation, written in Matlab, is publicly available [33]. It does not support distributed memory parallelism, and is restricted to conforming meshes that can be mapped to a square (in 2D) or a cube (in 3D). While, in practice, matrix assembly for high-order discretizations is discouraged, we use sparse assembled operators in this prototype implementation. In order to obtain more accurate results, we oversampled the domain $\Omega$ where $2\times 2$ elements per constant piece of the optical potential. We need to emphasize that the optical potentials just refers to the coordinates of scatterers and do not take into account the radii of each scatterer. That is to say we treat the scattering medium to be a point distribution, which is not the case for the fabricated nanohole membranes. However, such approximation does not affect the accuracy of the capability of the method to predict eigenmodes because the fabricated nanoholes have very small radii comparing to the multiscale nature of the scatterer separation. The multifractal potential term is added to the discrete Laplacian operator to obtain the discrete Helmholtz operator $H$. Our numerical simulations were performed using the Boston University’s Shared Computing Cluster (SCC) [34].
Figure 5: (a-h) Eigenmodes whose spectral parameters vary in the range $0.80.55<\overline{d}_{1}/\lambda<$ 0.8. (i) Superposition of eigenmodes corresponding to a narrow range (within a $5\%$ dispersion value) of eigenvalues around $k_{0}^{2}$. Incident wavelengths $\lambda$ are indicated in each panel.
@@
In our simulations, we selected the material permittivity to be $\epsilon_{r}=4$, corresponding to the value for the silicon nitride (SiN) material used in the fabrication of our devices. We chose $\overline{d}_{1}/\lambda=0.55$ and $\overline{d}_{1}/\lambda=0.62$ to compute the landscape and the eigenmodes of monofractal and multifractal systems respectively. The selected $\overline{d}_{1}/\lambda$ lies in the typical range for a photonic crystal structure where bandgaps are likely to form [31, 22] and optical modes with high-quality factors appear at the band-edges [31, 35, 36]. We emphasize that the range of the considered spectral parameter $\overline{d}_{1}/\lambda$ used in the simulations overlaps with the one of the fabricated structures that will be investigated in the experimental section of this paper.
Figure 6: (a-i) Eigenmodes whose spectral parameters vary in the range $0.780.6<\overline{d}_{1}/\lambda<$ 0.78. (i) Superposition of eigenmodes corresponding to a narrow range (within a $5\%$ dispersion value) of eigenvalues around $k_{0}^{2}$. Incident wavelengths $\lambda$ are indicated in each panel.
@@
The landscape function $u(\mathbf{r})$ of both monofractal and multifractal structures are plotted in Figures 4(c,d). One can see that the landscape of the monofractal potential features intense maxima localized within small groups of locally-symmetric clusters of adjacent scattering particles distributed across the structure with spatial distributions that closely resemble its fractal support. These are the regions in which the localization landscape predicts the existence of highly confined resonant modes localized by proximity effects occurring among closely coupled particles. This is the typical behavior observed in small fractal aggregates of dipolar particles [37]. On the other hand, the landscape of the multifractal potential displays localization regions that are more broadly distributed across the support of the structure. This is consistent with the non-homogeneous multiscale nature of multifractal systems [20]. In Figures 5(a-h) and 6(a-h), we display the calculated eigenmodes at the wavelengths specified in each panel. As for the 1D simulations, particular care has been taken to ensure that the eigenvectors are narrowly distributed around $k_{0}^{2}$ (within a $5\%$ dispersion value). Moreover, in order to facilitate comparisons with the experimental measurements in which multiple modes are simultaneously excited by an external focused beam, we also plotted in Figures 5(i) and 6(i) a linear combination of the modes shown in panels (a-h). As predicted by the localization landscape, we obtain strongly localized modes within small particle clusters for the monofractal structure, which are typical of self-similar arrays where they have been successfully exploited to dramatically increase the cross sections of nonlinear and Raman processes [37]. However, the situation is quite different in the multifractal structure, where high-intensity localized modes appear to spread across a more extended region of the geometrical support, reflecting its larger degree of spatial non-uniformity compared to the ones of monofractals. This is evident by looking at the mode superposition in Figure 6(i) in which the amplitude peaks cover a larger fraction of the structure. In the next sections, we discuss the fabrication and characterization of dielectric membrane structures designed based on the localization landscape theory and we demonstrate experimentally the formation of characteristic fractal modes in qualitative agreement with the landscape predictions.
@@ 5.1 Silicon nitride thin film growth and optical characterization
In this section we discuss the optical characterization of highly transparent silicon nitride thin films grown by reactive magnetron sputtering. Silicon nitride has become a well established material platform for integrated Si photonics where devices such as ultralow-loss waveguides [38], high-Q resonators, and integrated structures for telecommunications, sensing, and metrology have been demonstrated [39]. In this work, silicon nitride thin films were grown atop of silicon and fused silica substrates via reactive RF magnetron sputtering using a Denton Discovery 18 sputtering system with a base pressure of $2\times 10^{-7}$ Torr and a substrate temperature held at $300^{\circ}$. We used a 99.99% purity 3-in silicon target and sputtered the thin films in a 2:1 argon-nitrogen environment at 2.5 mTorr deposition pressure and 200 W of RF power resulting in a deposition rate of $\approx 5\,\mathrm{nm/min}$. All substrates were washed in a hot piranha solution and plasma ashed in an oxygen environment prior to deposition. We characterize the optical constants of the fabricated SiN thin films from the ultra-violet (UV) to the near infrared (NIR) via broadband variable angle spectroscopic ellipsometry (VASE) and normal incidence transmission measurements. In spectroscopic ellipsometry, a parameterized oscillator model is introduced to capture the relevant aspects of a materials optical dispersion as a function of wavelength. The oscillator model is then used to interpret the measured data and extract the optical constants of materials via accurate fitting procedures [40]. VASE measurements at three angles (${}^{\circ}$${}^{\circ}$${}^{\circ}$) were performed on SiN thin films grown atop of silicon substrates along with normal incidence transmission measurements of the same films grown atop of fused silica substrates. Figure 7 summarizes our work on characterizing the optical properties of the fabricated SiN thin films. Panels (a-b) show the measured and fitted ellipsometric parameters $\Psi$ and $\Delta$ corresponding to the magnitude and phase of the ratio of the complex s- and p-polarized Fresnel reflection coefficients, respectively, at each angle along with each respective fit. The complex Fresnel reflection coefficients are related to the complex refractive index and thickness of the material through parameterized oscillator models. In particular, we employ a single parameterized Tauc-Lorentz (TL) oscillator model which has been used to extract the optical constants of amorphous materials in excellent agreement with experimental measurements [41]. We find that this modeling method gives nearly ideal agreement with our measurements over the entire UV-VIS-NIR region investigated with a mean square error (MSE) of less than 10. In Figure 7(c) we report the extracted optical constants of the fabricated 350 nm thick SiN thin film as a function of wavelength. Furthermore, we performed normal incidence transmission measurements on the SiN thin films grown atop of fused silica substrates and independently test our TL model by plotting the expected transmission. We report our measurements and fit in fig 7(d) finding the TL model is in excellent agreement with measurement over the entire wavelength range investigated. The transmission spectra, along with the extracted optical constants of the fabricated material, demonstrates exceptionally high transparency with an almost constant refractive index of $2$. Consistent with literature, we attribute these optical properties to the high RF power and substrate temperature during deposition[42]. In the next section we discuss the fabrication of nanohole perforated membranes with multifractal geometries.